OpenStudy (anonymous):

if a number is increased by 12 and then divided by 3, the result is 7

OpenStudy (secretx_x):

Okay?

OpenStudy (secretx_x):

Do you need some form of help? Cause this isn't very descriptive.

jimthompson5910 (jim_thompson5910):

"a number is increased by 12 and then divided by 3, the result is 7" translates to the equation \[\Large \frac{x+12}{3} = 7\] where x is the unknown number

jimthompson5910 (jim_thompson5910):

Multiply both sides by 3 to get \[\Large \frac{x+12}{3} = 7\] \[\Large 3*\frac{x+12}{3} = 3*7\] \[\Large \cancel{3}*\frac{x+12}{\cancel{3}} = 3*7\] \[\Large x+12 = 3*7\] \[\Large x+12 = 21\] what's the next step?